# Bilinear spherical maximal function with fractal dilations

### v1

@@ 5. Estimate for local maximal operator
Let $0\leq\beta\leq 1$. Then for any $\mathbf{n}\in{\mathbb{N}}^{2}$, we have
@@
    $\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\|_{L^{1}}\lesssim 2^{-|\mathbf{n}|\frac{2d-2-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}},$
@@
    $\left\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\right\|_{L^{2/3}}\lesssim 2^{-|\mathbf{n}|\frac{d-3-\beta}{2}}\|f_{1}\|_{L^{1}}\|f_{2}\|_{L^{2}},$
  3. (3)
    $\left\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\right\|_{L^{2/3}}\lesssim 2^{-|\mathbf{n}|\frac{d-3-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{1}}.$
@@
Let $0\leq\beta\leq 1$. Then for any $\mathbf{n}\in{\mathbb{N}}^{2}$, we have
@@
Note that if we assume that Lemma 5.2 and Lemma 5.3 hold true, then bilinear analogue of real interpolation theorem, see Theorem 7.2.2 in [modernGrafakos], yields strong-type $L^{p}$-estimates
@@
for some $\epsilon>0$ and $\big(\frac{1}{p_{1}},\frac{1}{p_{2}}\big)\in\Omega(\{O,A,Q,P,R,B\})\setminus\{A,B,PQ,PR\}$. Therefore, summing over $n_{1}$ and $n_{2},$ we get the desired strong-type $L^{p}$-estimates for the operator $\mathcal{M}_{E_{0}}$.
Moreover, the bilinear complex interpolation theorem applied to estimates $(1)$ and $(2)$ of Lemma 5.2 gives us
$\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\|_{L^{\frac{2p}{p+2}}}\lesssim 2^{-|\mathbf{n}|\big(\frac{3d-1-\beta}{2}-\frac{d+1}{p}\big)}\|f_{1}\|_{L^{p}}\|f_{2}\|_{L^{2}},\quad 1\leq p\leq 2.$
Finally, for the operator $\mathcal{M}_{E_{0}}$, the restricted weak-type inequality at the point $\big(\frac{1}{p_{1}},\frac{1}{p_{2}}\big)$ lying on the line segment $PQ$ follows by Bourgain’s interpolation Lemma 5.1, applied to the preceding estimate and Lemma 5.3. The same argument could be used to obtain restricted weak-type estimates for points on the line segment $PR$. ∎
Therefore, we are left with proving Lemma 5.2 and Lemma 5.3.
@@
Let $\mathcal{E}_{0}$ be the disjoint collection of intervals of length $2^{-|\mathbf{n}|}$ covering $E_{0}$. Then, we can write
$\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})(x)$ $=\sup_{I\in\mathcal{E}_{0}}\sup_{t\in I}\Bigg|\mathcal{A}_{t}\Big(R_{n_{1}}\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}F_{N}f_{1},R_{n_{2}}f_{2}\Big)(x)\Bigg|,$
where $\widehat{F_{N}(f)}=\chi_{{}_{N}}\widehat{f}$.
For $t\in I$, using the Fundamental Theorem of Calculus we get that
$\sup_{t\in I}\Big|\mathcal{A}_{t}(R_{n_{1}}F_{N}f_{1},R_{n_{2}}f_{2})(x)\Big|$
$\lesssim$ $\left(\frac{1}{|I|}\int_{I}|\mathcal{A}_{s}(R_{n_{1}}F_{N}f_{1},R_{n_{2}}f_{2})(x)|^{2}\;ds\right)^{\frac{1}{2}}$
$+\left(\int_{I}\mathcal{A}_{s}(R_{n_{1}}F_{N}f_{1},R_{n_{2}}f_{2})(x)\frac{\partial}{\partial s}\mathcal{A}_{s}(R_{n_{1}}F_{N}f_{1},R_{n_{2}}f_{2})(x)\;ds\right)^{\frac{1}{2}}$
$\leq$ $\left(\frac{1}{|I|}\int_{I}|\mathcal{A}_{s}(R_{n_{1}}f_{1},R_{n_{2}}f_{2})(x)|^{2}\;ds\right)^{\frac{1}{2}}$
$+\left(\int_{I}|\mathcal{A}_{s}(R_{n_{1}}F_{N}f_{1},R_{n_{2}}f_{2})(x)|^{2}\;ds\right)^{\frac{1}{4}}\left(\int_{I}\left|\frac{\partial}{\partial s}\mathcal{A}_{s}(R_{n_{1}}F_{N}f_{1},R_{n_{2}}f_{2})(x)\right|^{2}\;ds\right)^{\frac{1}{4}}.$
Using $\ell^{2}\hookrightarrow\ell^{\infty}$ embedding, we get that
$\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})(x)\lesssim$ $\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}2^{\frac{|\mathbf{n}|}{2}}\mathcal{G}^{\mathbf{n}}_{0,N}(f_{1},f_{2})(x)+\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\big(\mathcal{G}^{\mathbf{n}}_{0,N}(f_{1},f_{2})(x)\tilde{\mathcal{G}}^{\mathbf{n}}_{0,N}(f_{1},f_{2})(x)\big)^{\frac{1}{2}},$ (5.1)
@@
$\mathcal{G}^{\mathbf{n}}_{0,N}(f_{1},f_{2})(x):=\left(\sum_{I\in\mathcal{E}_{0}}\int_{I}|\mathcal{A}_{s}(R_{n_{1}}F_{N}f_{1},R_{n_{2}}f_{2})(x)|^{2}\;ds\right)^{\frac{1}{2}}~\text{ and}$
$\tilde{\mathcal{G}}^{\mathbf{n}}_{0,N}(f_{1},f_{2})(x):=\left(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|\frac{\partial}{\partial s}\mathcal{A}_{s}(R_{n_{1}}F_{N}f_{1},R_{n_{2}}f_{2})(x)\right|^{2}\;ds\right)^{\frac{1}{2}}.$
@@
$\left\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\right\|_{L^{p}}$
$\lesssim 2^{\frac{|\mathbf{n}|}{2}}\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\mathcal{G}^{\mathbf{n}}_{0,N}(f_{1},f_{2})\Bigg\|_{L^{p}}+\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\mathcal{G}^{\mathbf{n}}_{0,N}(f_{1},f_{2})\Bigg\|_{L^{p}}^{\frac{1}{2}}\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\tilde{\mathcal{G}}^{\mathbf{n}}_{0,N}(f_{1},f_{2})\Bigg\|_{L^{p}}^{\frac{1}{2}}.$ (5.2)
@@
Note that the multiplier $2^{-|\mathbf{n}|}\frac{\partial}{\partial s}\widehat{\sigma}(s\xi,s\eta)$ behaves like $\widehat{\sigma}(s\xi,s\eta)$ where $1\leq s\leq 2$. Therefore, it is enough to prove the desired estimates for the square function $\Big(\sum\limits_{I\in\mathcal{E}_{0}}\int_{I}|\mathcal{A}_{s}(R_{n_{1}}F_{N}f_{1},R_{n_{2}}f_{2})(x)|^{2}\;ds\Big)^{\frac{1}{2}}.$ We invoke the decomposition of the multiplier carried out in Section 4 and use the estimate (4.9) to get that
$\mathcal{G}^{\mathbf{n}}_{0,N}(f_{1},f_{2})(x)$
@@
$\lesssim\left(\int_{\mathbb{R}^{d}}\sup_{1\leq s\leq 2}\left|B^{\mathbf{n}}_{s,N}\left(F_{N}f_{1}\right)(x,\eta)\right|d\eta\right)\bigg(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|L_{n_{2}}\left(U_{s,N}^{n_{2}}f_{2}\right)(x)\right|^{2}\;ds\bigg)^{\frac{1}{2}}+error~terms.$
Note that for the error terms, we have better decay in the multiplier, see (4.2). Thus, it is enough to prove $L^{p}$-estimates for the following operator
@@
$f_{1}\rightarrow\left(\int_{\mathbb{R}^{d}}\sup_{1\leq s\leq 2}\left|B^{\mathbf{n}}_{s,N}\left(F_{N}f_{1}\right)(x,\eta)\right|d\eta\right)$

### v2

@@ 5. Estimate for local maximal operator $\mathcal{M}_{E_{0}}$
Let $d\geq 1$ and $0\leq\beta\leq 1$. Then for any $\mathbf{n}\in{\mathbb{N}}^{2}$, we have
$\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\|_{L^{1}}\lesssim 2^{-|\mathbf{n}|\frac{2d-1-2\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}.$
Lemma 5.2.
Let $d\geq 3$ and $0\leq\beta\leq 1$. Then for any $\mathbf{n}\in{\mathbb{N}}^{2}$, we have
@@
    $\left\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\right\|_{L^{2/3}}\lesssim\min\left\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{2}},2^{-|\mathbf{n}|\frac{d-3-\beta}{2}}\right\}\|f_{1}\|_{L^{1}}\|f_{2}\|_{L^{2}},$
@@
    $\left\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\right\|_{L^{2/3}}\lesssim\min\left\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{2}},2^{-|\mathbf{n}|\frac{d-3-\beta}{2}}\right\}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{1}}.$
@@
Let $d\geq 1$ and $0\leq\beta\leq 1$. Then for any $\mathbf{n}\in{\mathbb{N}}^{2}$, we have
@@
Remark 5.1.
We would like to remark that by an application of Bourgain’s interpolation argument [[30, Lemma 2.6]] along with the estimates established in Lemma 5.1, Lemma 5.2, and Lemma 5.3, we get that for all measurable sets $F_{1}$ and $F_{2}$, the following estimate holds
$\mathcal{M}_{E_{0}}(\chi_{F_{1}},\chi_{F_{2}})||_{L^{p,\infty}}\lesssim|F_{1}|^{\frac{1}{p_{1}}}|F_{2}|^{\frac{1}{p_{2}}}$
holds whenever the point $(\frac{1}{p_{1}},\frac{1}{p_{2}})$ lies on the line segment $PQ$ or $PR$ in dimensions $d\geq 3$. However, observe that for $p<1$, the estimate above does not allow us to deduce restricted weak-type bounds for the local maximal function $\mathcal{M}_{E_{0}}(f_{1},f_{2}).$
Note that if we assume that Lemma 5.1, Lemma 5.2 and Lemma 5.3 hold true, then the bilinear analogue of real interpolation theorem, see Theorem 7.2.2 in [23], yields strong-type $L^{p}$-estimates
@@
for some $\epsilon>0$ and $\big(\frac{1}{p_{1}},\frac{1}{p_{2}}\big)\in\Omega(O,A,Q,P,R,B)\setminus\{A,B,PQ,PR\}$. Therefore, summing over $n_{1}$ and $n_{2},$ we get the desired strong-type $L^{p}$-estimates for the operator $\mathcal{M}_{E_{0}}$.
Therefore, we are left with proving Lemma 5.1, Lemma 5.2 and Lemma 5.3.
@@
From (4.1) and the asymptotics of the Bessel function, we have the following decomposition of $\mathcal{M}^{\mathbf{n}}_{0}$.
$\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})(x)\leq\mathbb{I}^{\mathbf{n}}(x)+\mathbb{II}^{\mathbf{n}}(x)+\mathbb{III}^{\mathbf{n}}(x),$ (5.1)
@@
$\mathbb{I}^{\mathbf{n}}(x)=\sup_{t\in E_{0}}\left|\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\frac{a_{0}e^{it|(\xi,\eta)|}}{(t|(\xi,\eta)|)^{\frac{2d-1}{2}}}\widehat{\psi}\left(2^{-n_{1}}\xi\right)\widehat{\psi}\left(2^{-n_{2}}\eta\right)\widehat{f_{1}}(\xi)\widehat{f_{2}}(\eta)e^{ix\cdot(\xi+\eta)}~d\xi d\eta\right|,$
$\mathbb{II}^{\mathbf{n}}(x)=\sup_{t\in E_{0}}\left|\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\frac{b_{0}e^{-it|(\xi,\eta)|}}{(t|(\xi,\eta)|)^{\frac{2d-1}{2}}}\widehat{\psi}\left(2^{-n_{1}}\xi\right)\widehat{\psi}\left(2^{-n_{2}}\eta\right)\widehat{f_{1}}(\xi)\widehat{f_{2}}(\eta)e^{ix\cdot(\xi+\eta)}~d\xi d\eta\right|,$
and
$\mathbb{III}^{\mathbf{n}}(x)=\sup_{t\in E_{0}}\left|\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\mathfrak{m}_{t}^{\mathbf{n}}(\xi,\eta)\widehat{f_{1}}(\xi)\widehat{f_{2}}(\eta)e^{ix\cdot(\xi+\eta)}~d\xi d\eta\right|.$
Here $\mathfrak{m}_{t}^{\mathbf{n}}(\xi,\eta)=\Big[\widehat{d\sigma}(t\xi,t\eta)-\Big(\frac{a_{0}e^{it|(\xi,\eta)|}}{(t|(\xi,\eta)|)^{\frac{2d-1}{2}}}+\frac{b_{0}e^{-it|(\xi,\eta)|}}{(t|(\xi,\eta)|)^{\frac{2d-1}{2}}}\Big)\Big]\widehat{\psi}\left(2^{-n_{1}}\xi\right)\widehat{\psi}\left(2^{-n_{2}}\eta\right)$. Using the estimate (4.2) with $M=0$, we have
$\big|\partial^{\alpha}_{\xi}\partial^{\beta}_{\eta}\mathfrak{m}_{t}^{\mathbf{n}}(\xi,\eta)\big|\lesssim 2^{-|\mathbf{n}|\frac{2d-1}{2}}|\xi|^{-|\alpha|}|\eta|^{-|\beta|},\text{ uniformly in }t\in E_{0}.$
Since the multiplier $\mathfrak{m}_{t}^{\mathbf{n}}$ is compactly supported away from origin and satisfies the above derivative condition, we get that
$\mathbb{III}^{\mathbf{n}}(x)\lesssim 2^{-|\mathbf{n}|\frac{2d-1}{2}}M_{HL}f_{1}(x)M_{HL}f_{2}(x).$ (5.2)
Therefore, $\mathbb{III}^{\mathbf{n}}$ satisfies the desired bounds in Lemma 5.1. It remains to prove the estimates in Lemma 5.1 for $\mathbb{I}^{\mathbf{n}}$ and $\mathbb{II}^{\mathbf{n}}$. We will establish the bounds for $\mathbb{I}^{\mathbf{n}}$, and the same argument can be applied to $\mathbb{II}^{\mathbf{n}}$ to obtain identical bounds.
Let $\mathcal{E}_{0}$ be the collection of disjoint intervals of length $2^{-|\mathbf{n}|}$ covering $E_{0}$. Then, we can write
$\mathbb{I}^{\mathbf{n}}(x)$ $=\sup_{t\in E_{0}}\left|\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\frac{a_{0}e^{it|(\xi,\eta)|}}{(t|(\xi,\eta)|)^{\frac{2d-1}{2}}}\widehat{\psi}\left(2^{-n_{1}}\xi\right)\widehat{\psi}\left(2^{-n_{2}}\eta\right)\Bigg(\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\widehat{F_{N}f_{1}}(\xi)\Bigg)\widehat{f_{2}}(\eta)e^{ix\cdot(\xi+\eta)}~d\xi d\eta\right|$
$\leq|a_{0}|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\sup_{I\in\mathcal{E}_{0}}\sup_{t\in I}\Bigg|S^{\mathbf{n}}_{t,N}(f_{1},f_{2})(x)\Bigg|,$
where the operators $F_{N}$ and $S^{\mathbf{n}}_{t,N}$ are defined in (4.4) and (4.5) respectively.
For $t\in I$, using the Fundamental Theorem of Calculus, we get that
$\sup_{t\in I}\Big|S^{\mathbf{n}}_{t,N}(f_{1},f_{2})(x)\Big|$ (5.3)
$\lesssim$ $\left(\frac{1}{|I|}\int_{I}|S^{\mathbf{n}}_{s,N}(f_{1},f_{2})(x)|^{2}\;ds\right)^{\frac{1}{2}}+\left(\int_{I}S^{\mathbf{n}}_{s,N}(f_{1},f_{2})(x)\frac{\partial}{\partial s}S^{\mathbf{n}}_{s,N}(f_{1},f_{2})(x)\;ds\right)^{\frac{1}{2}}$
$\leq$ $\left(\frac{1}{|I|}\int_{I}|S^{\mathbf{n}}_{s,N}(f_{1},f_{2})(x)|^{2}\;ds\right)^{\frac{1}{2}}+\left(\int_{I}|S^{\mathbf{n}}_{s,N}(f_{1},f_{2})(x)|^{2}\;ds\right)^{\frac{1}{4}}\left(\int_{I}\left|\frac{\partial}{\partial s}S^{\mathbf{n}}_{s,N}(f_{1},f_{2})(x)\right|^{2}\;ds\right)^{\frac{1}{4}}.$
Using $\ell^{1}\hookrightarrow\ell^{\infty}$ embedding for $\mathcal{E}_{0}$, we get that
$\mathbb{I}^{\mathbf{n}}(x)\lesssim$ $\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\sum_{I\in\mathcal{E}_{0}}2^{\frac{|\mathbf{n}|}{2}}\mathcal{G}^{\mathbf{n}}_{I,N}(f_{1},f_{2})(x)+\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\sum_{I\in\mathcal{E}_{0}}\big(\mathcal{G}^{\mathbf{n}}_{I,N}(f_{1},f_{2})(x)\tilde{\mathcal{G}}^{\mathbf{n}}_{I,N}(f_{1},f_{2})(x)\big)^{\frac{1}{2}},$
where
$\mathcal{G}^{\mathbf{n}}_{I,N}(f_{1},f_{2})(x):=\left(\int_{I}|S^{\mathbf{n}}_{s,N}(f_{1},f_{2})(x)|^{2}\;ds\right)^{\frac{1}{2}}~\text{ and}$
$\tilde{\mathcal{G}}^{\mathbf{n}}_{I,N}(f_{1},f_{2})(x):=\left(\int_{I}\left|\frac{\partial}{\partial s}S^{\mathbf{n}}_{s,N}(f_{1},f_{2})(x)\right|^{2}\;ds\right)^{\frac{1}{2}}.$
@@
$\left\|\mathbb{I}^{\mathbf{n}}\right\|_{L^{p}}$ (5.4)
$\lesssim 2^{\frac{|\mathbf{n}|}{2}}\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\sum_{I\in\mathcal{E}_{0}}\mathcal{G}^{\mathbf{n}}_{I,N}(f_{1},f_{2})\Bigg\|_{L^{p}}+\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\sum_{I\in\mathcal{E}_{0}}\mathcal{G}^{\mathbf{n}}_{I,N}(f_{1},f_{2})\Bigg\|_{L^{p}}^{\frac{1}{2}}\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\sum_{I\in\mathcal{E}_{0}}\tilde{\mathcal{G}}^{\mathbf{n}}_{I,N}(f_{1},f_{2})\Bigg\|_{L^{p}}^{\frac{1}{2}}.$
@@
Since the multiplier $2^{-|\mathbf{n}|}\frac{\partial}{\partial s}\frac{e^{is|(\xi,\eta)|}}{(s|(\xi,\eta)|)^{\frac{2d-1}{2}}}$ behaves like $\frac{e^{is|(\xi,\eta)|}}{(s|(\xi,\eta)|)^{\frac{2d-1}{2}}}$ for $1\leq s\leq 2$, it is enough to prove $L^{p}$-estimates for the square function $\Big(\int_{I}|S^{\mathbf{n}}_{s,N}(f_{1},f_{2})(x)|^{2}\;ds\Big)^{\frac{1}{2}}.$ We invoke the decomposition of the operator $S^{\mathbf{n}}_{t,N}$ from (4.11) to get that
@@
$\mathcal{G}^{\mathbf{n}}_{I,N}(f_{1},f_{2})(x)\lesssim\left(\int_{\mathbb{R}^{d}}\sup_{s\in I}\left|B^{\mathbf{n}}_{s,N}\left(F_{N}f_{1}\right)(x,\eta)\right|d\eta\right)\bigg(\int_{I}\left|L_{n_{2}}\left(U_{s,N}^{n_{2}}f_{2}\right)(x)\right|^{2}\;ds\bigg)^{\frac{1}{2}}.$
Therefore, it is enough to prove $L^{p}$-estimates for the following operators
$f_{1}\rightarrow\int_{\mathbb{R}^{d}}\sup_{s\in I}\left|B^{\mathbf{n}}_{s,N}\left(F_{N}f_{1}\right)(x,\eta)\right|d\eta$
and
$f_{2}\rightarrow\bigg(\int_{I}\left|L_{n_{2}}\left(U_{s,N}^{n_{2}}f_{2}\right)(x)\right|^{2}\;ds\bigg)^{\frac{1}{2}}.$
Observe that using Minkowski’s integral inequality, we have
$\bigg(\int_{I}\left|L_{n_{2}}\left(U_{s,N}^{n_{2}}f_{2}\right)(x)\right|^{2}\;ds\bigg)^{\frac{1}{2}}\leq$ $\int_{{\mathbb{R}}^{d}}\bigg(\int_{I}\left|U_{s,N}^{n_{2}}f_{2}(x-y)\right|^{2}\;ds\bigg)^{\frac{1}{2}}\frac{dy}{\left(1+4\pi^{2}\left|2^{n_{2}}y\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}.$ (5.5)
We take $L^{2}$-norm in the above and use Plancherel’s theorem to get that
$\left\|\bigg(\int_{I}\left|L_{n_{2}}\left(U_{s,N}^{n_{2}}f_{2}\right)\right|^{2}\;ds\bigg)^{\frac{1}{2}}\right\|_{L^{2}}\leq$ $\left\|\bigg(\int_{I}\big|U_{s,N}^{n_{2}}f_{2}\big|^{2}\;ds\bigg)^{\frac{1}{2}}\right\|_{L^{2}}\int_{{\mathbb{R}}^{d}}\frac{dy}{\left(1+4\pi^{2}\left|2^{n_{2}}y\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}$ (5.6)
$\lesssim$ $2^{-n_{2}d}2^{-\frac{|\mathbf{n}|}{2}}\|f_{2}\|_{L^{2}}.$
Next, by Minkowski’s integral inequality and an application of Lemma 4.1, we obtain
$\left\|\int_{\mathbb{R}^{d}}\sup_{s\in I}\left|B^{\mathbf{n}}_{s,N}\left(F_{N}f_{1}\right)(\cdot,\eta)\right|d\eta\right\|_{L^{2}}\lesssim 2^{-|\mathbf{n}|d}2^{n_{2}d}\|F_{N}f\|_{L^{2}}.$ (5.7)
Using Cauchy Schwarz inequality along with the estimates (5.6) and (5.7), we obtain
$\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\sum_{I\in\mathcal{E}_{0}}\mathcal{G}^{\mathbf{n}}_{I,N}(f_{1},f_{2})\Bigg\|_{L^{1}}\lesssim$ $2^{-|\mathbf{n}|\frac{2d+1}{2}}\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\sum_{I\in\mathcal{E}_{0}}\|F_{N}f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}$
$\lesssim$ $2^{-|\mathbf{n}|\frac{2d-2\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}},$
where we used the cardinality bound $\mathcal{E}_{0}\leq 2^{|\mathbf{n}|\beta}$, which is a direct consequence of $E$ having Minkowski dimension $\beta$ (see Definition 1.1).
As described earlier, similar arguments as above, yield the following bounds for the square function involving the derivative term $\frac{\partial}{\partial s}S^{\mathbf{n}}_{t,N}$.
$\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\sum_{I\in\mathcal{E}_{0}}\tilde{\mathcal{G}}^{\mathbf{n}}_{I,N}(f_{1},f_{2})\Bigg\|_{L^{1}}\lesssim 2^{-|\mathbf{n}|\frac{2d-2-2\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}},$
Substituting the estimates obtained above in (5.4), we have
$\left\|\mathbb{I}^{\mathbf{n}}\right\|_{L^{1}}\lesssim$ $2^{\frac{|\mathbf{n}|}{2}}2^{-|\mathbf{n}|\frac{2d-2\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}+\left(2^{-|\mathbf{n}|\frac{2d-2\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}\right)^{\frac{1}{2}}\left(2^{-|\mathbf{n}|\frac{2d-2-2\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}\right)^{\frac{1}{2}}$
$\lesssim$ $2^{-|\mathbf{n}|\frac{2d-1-2\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}.$
This proves the desired estimates for $\mathbb{I}^{\mathbf{n}}$, and similarly for $\mathbb{II}^{\mathbf{n}}$. These estimates, together with (5.2), complete the proof of Lemma 5.2. ∎
5.2. Proof of Lemma 5.2.
Based on the decomposition of $\mathcal{M}^{\mathbf{n}}_{0}$ given in (5.1) and the estimates for $\mathbb{III}^{\mathbf{n}}$ in (5.2), it is enough to prove the desired estimates in Lemma 5.2 for $\mathbb{I}^{\mathbf{n}}$ and $\mathbb{II}^{\mathbf{n}}$. Again, we prove the bounds for $\mathbb{I}^{\mathbf{n}}$, and the same bounds follow for $\mathbb{II}^{\mathbf{n}}$ by a similar argument.
The estimate from (5.5) along with Lemma 4.2 gives us
$\left\|\bigg(\int_{I}\left|L_{n_{2}}\left(U_{s,N}^{n_{2}}f_{2}\right)\right|^{2}\;ds\bigg)^{\frac{1}{2}}\right\|_{L^{1}}\lesssim 2^{-n_{2}\frac{d}{2}}2^{-\frac{|\mathbf{n}|}{2}}\|f_{2}\|_{L^{1}}.$ (5.8)
Now, applying Hölder’s inequality along with the estimates (5.7) and (5.8), we obtain
$\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\sum_{I\in\mathcal{E}_{0}}\mathcal{G}^{\mathbf{n}}_{I,N}(f_{1},f_{2})\Bigg\|_{L^{2/3}}$ $\lesssim 2^{-|\mathbf{n}|\frac{d+1}{2}}\bigg(\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\sum_{I\in\mathcal{E}_{0}}\|F_{N}f_{1}\|_{L^{2}}^{\frac{2}{3}}\|f_{2}\|_{L^{1}}^{\frac{2}{3}}\bigg)^{\frac{3}{2}}$
$\lesssim 2^{-|\mathbf{n}|\frac{d-1-3\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{1}},,$
and
$\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\sum_{I\in\mathcal{E}_{0}}\tilde{\mathcal{G}}^{\mathbf{n}}_{I,N}(f_{1},f_{2})\Bigg\|_{L^{2/3}}$ $\lesssim 2^{-|\mathbf{n}|\frac{d-3-3\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{1}}.$
Combining the estimates for square functions together with equation (5.4), we obtain
$\left\|\mathbb{I}^{\mathbf{n}}\right\|_{L^{2/3}}\lesssim 2^{-|\mathbf{n}|\frac{d-2-3\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{1}}.$ (5.9)
This proves one of the bounds in Lemma 5.2 for $\mathbb{I}^{\mathbf{n}}$. For the other bound, the $\ell^{2}\hookrightarrow\ell^{\infty}$ embedding for $\mathcal{E}_{0}$ along with (5.3) gives us
$\mathbb{I}^{\mathbf{n}}(x)\lesssim$ $\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}2^{\frac{|\mathbf{n}|}{2}}\mathcal{H}^{\mathbf{n}}_{0,N}(f_{1},f_{2})(x)+\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\big(\mathcal{H}^{\mathbf{n}}_{0,N}(f_{1},f_{2})(x)\tilde{\mathcal{H}}^{\mathbf{n}}_{0,N}(f_{1},f_{2})(x)\big)^{\frac{1}{2}},$ (5.10)
where
$\mathcal{H}^{\mathbf{n}}_{0,N}(f_{1},f_{2})(x):=\left(\sum_{I\in\mathcal{E}_{0}}\int_{I}|S^{\mathbf{n}}_{s,N}(f_{1},f_{2})(x)|^{2}\;ds\right)^{\frac{1}{2}}~\text{ and}$
$\tilde{\mathcal{H}}^{\mathbf{n}}_{0,N}(f_{1},f_{2})(x):=\left(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|\frac{\partial}{\partial s}S^{\mathbf{n}}_{s,N}(f_{1},f_{2})(x)\right|^{2}\;ds\right)^{\frac{1}{2}}.$
Using the decomposition of the operator $S^{\mathbf{n}}_{t,N}$ from (4.11), we get that
@@
$\mathcal{H}^{\mathbf{n}}_{0,N}(f_{1},f_{2})(x)\lesssim\left(\int_{\mathbb{R}^{d}}\sup_{1\leq s\leq 2}\left|B^{\mathbf{n}}_{s,N}\left(F_{N}f_{1}\right)(x,\eta)\right|d\eta\right)\bigg(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|L_{n_{2}}\left(U_{s,N}^{n_{2}}f_{2}\right)(x)\right|^{2}\;ds\bigg)^{\frac{1}{2}}.$
Therefore, it is enough to prove $L^{p}$-estimates for the following operators
$f_{1}\rightarrow\int_{\mathbb{R}^{d}}\sup_{1\leq s\leq 2}\left|B^{\mathbf{n}}_{s,N}\left(F_{N}f_{1}\right)(x,\eta)\right|d\eta$